A stero cost $677.25 after 7.5% sales tax. Find the price before tax.
I'm having issues with converting the sales tax to a decimal so that I can divide it into the total cost in order to find the price before taxes were applied in the form of a operational formula.

Answer by checkley77(12844)

X+.075X=677.25
1.075X=677.25
X=677.25/1.075
X=$630.00 COST OF THE STERO BEFORE TAXES.
PROOF:
630.00+.075*630.00=677.25
630.00+47.25=677.25
677.25=677.25